Prostaglandins are a class of eicosanoids, fatty acid derivatives with a variety of extremely potent actions on vertebrate tissues. They are responsible for producing fever and inflammation and its associated pain. Prostaglandins are derived from the 20-carbon fatty acid arachidonic acid in a reaction catalyzed by the enzyme prostaglandin endoperoxide synthase. This enzyme, a cyclooxygenase, uses oxygen to convert arachidonic acid to PGG2, the immediate precursor of many different prostaglandins.

Arachidonic acid (mM) Rate of formation of PGG2 (mM/min) Rate of formation of PGG2 with 10 mg/ml ibuprofen (mM/min)
0.275 17.3 10.8
0.330 19.5 12.5
0.495 24.6 16.7
0.825 31.2 22.9
1.93 40.4 33.7

The kinetic data given in the table are for the reaction catalyzed by a mutant of prostaglandin endoperoxide synthase. Focusing here on the first two columns, determine the ?max and ?m of the enzyme.

Find Vmax and Km.
Answer #1

Michaelis Menten equation is

v = Vmax / (1 + (Km/[S]))

Lineweaver burk plot is between 1/[S] on x-axis and 1/Vo on y-axis.

0.275 substrate(MM) velocity (MM/min) 1/ S 1 17.3 3.636364 0.33 19.5 3.030303 0.495 24.6 2.020202 0.825 31.2 1.212121 1.93 40

Equation of this graph is y = 0.0106x + 0.0192

According to this equation, slope = 0.0106 and intercept = 0.0192

Also, intercept = 1/Vmax = 0.0192

Vmax = 52.08 mM/min

Slope = Km/Vmax = 0.0106

So, Km = 0.0106 *Vmax = 0.0106 *52.08 = 0.551 mM

Answer : Km = 0.551 mM

Vmax = 52.08 mM/min